### v1

# The LSCD Benchmark: a Testbed for Diachronic Word Meaning Tasks

Abstract
Lexical Semantic Change Detection (LSCD) is a complex, lemma-level task, which is usually operationalized based on two subsequently applied usage-level tasks: First, Word-in-Context (WiC) labels are derived for pairs of usages. Then, these labels are represented in a graph on which Word Sense Induction (WSI) is applied to derive sense clusters. Finally, LSCD labels are derived by comparing sense clusters over time. This modularity is reflected in most LSCD datasets and models. It also leads to a large heterogeneity in modeling options and task definitions, which is exacerbated by a variety of dataset versions, preprocessing options and evaluation metrics. This heterogeneity makes it difficult to evaluate models under comparable conditions, to choose optimal model combinations or to reproduce results. Hence, we provide a benchmark repository standardizing LSCD evaluation. Through transparent implementation results become easily reproducible and by standardization different components can be freely combined. The repository reflects the task’s modularity by allowing model evaluation for WiC, WSI and LSCD. This allows for careful evaluation of increasingly complex model components providing new ways of model optimization.
1 Introduction
Lexical Semantic Change Detection (LSCD) is a field of NLP that studies methods automating the analysis of changes in word meanings over time. In recent years, this field has seen much development in terms of models, datasets and tasks (Schlechtweg, 2023). LSCD is a complex, lemma-level task, which is usually operationalized based on two subsequently applied usage-level tasks: First, Word-in-Context (WiC) labels are derived for pairs of usages. Then, these labels are represented in a graph on which Word Sense Induction (WSI) is applied to derive sense clusters. Finally, LSCD labels are derived by comparing sense clusters over time. This modularity is reflected in most LSCD datasets and models. It also leads to a large heterogeneity in modeling options and task definitions, which is exacerbated by a variety of dataset versions, preprocessing options and evaluation metrics. This heterogeneity makes it difficult to evaluate models under comparable conditions, to choose optimal model combinations or to reproduce results.
In order to handle this heterogeneity, we think that a shared testbed with a common evaluation setup is needed. Hence, we present a benchmark repository implementing evaluation procedures for models on most available LSCD datasets. [^1] https://github.com/ChangeIsKey/LSCDBenchmark The benchmark exploits the modularity of the meta task LSCD by allowing for evaluation of the subtasks WiC and WSI on the same datasets. It can be assumed that performance on the subtasks directly determines performance on the meta task. We aim to stimulate transfer between the fields of WiC, WSI and LSCD by providing a repository allowing for evaluation on all these tasks with shared model components.
We hope that the resulting benchmark by standardizing the evaluation of LSCD models and providing models with near-SOTA performance can serve as a starting point for researchers to develop and improve models. The benchmark allows for a wide application and testing of models by focusing on multilingual models and their evaluation on several languages.
2 Related Work
A number of recently created LSCD data sets apply WiC and WSI in the annotation process (cf. Section 3) and thus allow for evaluation of WiC and WSI along with LSCD models (i.a. Schlechtweg et al., 2021; Kurtyigit et al., 2021; Kutuzov et al., 2022; Zamora-Reina et al., 2022). [^2] The bulk of these data sets is listed at: https://www.ims.uni-stuttgart.de/data/wugs. There are also a number of data sets omitting WSI, but allowing for WiC and LSCD evaluation (i.a. Schlechtweg et al., 2018; Rodina and Kutuzov, 2020; Kutuzov and Pivovarova, 2021), or data sets omitting the WiC allowing for WSI and LSCD evaluation (i.a. Basile et al., 2020; Cook et al., 2014), or datasets allowing purely for WiC evaluation (Loureiro et al., 2022). Above that, there is a number of data sets allowing for synchronic evaluation of WiC (i.a. Pilehvar and Camacho-Collados, 2019; Trott and Bergen, 2021) or WSI (i.a. Langone et al., 2004; Hovy et al., 2006) or both (Erk et al., 2013; Aksenova et al., 2022), which can be exploited e.g. by simulating LSCD labels (Rosenfeld and Erk, 2018; Dubossarsky et al., 2019; Schlechtweg and Schulte im Walde, 2020). Note also that there is no restriction on the source or strategy for sampling word usages, i.e., they do not necessarily have to be sampled from different time periods, but could also be sampled from different text genres, domains, dialects or even languages. Hence, data sets reflecting meaning divergences between these text categories can be integrated into the benchmark as well (i.a. Hätty et al., 2019; Baldissin et al., 2022).
So far, there is no comprehensive LSCD benchmark, implementing state-of-the-art models on (human-annotated) high-quality evaluation data from multiple languages and multiple time periods. The leaderboards of several standalone shared tasks can be seen as small-scale benchmarks without common model implementation (Ahmad et al., 2020; Basile et al., 2020; Kutuzov and Pivovarova, 2021; Zamora-Reina et al., 2022) with the SemEval task being the most diverse with four languages (Schlechtweg et al., 2020). Schlechtweg et al. (2019) provide a comprehensive repository of type-based modeling approaches to LSCD with evaluation pipelines on multiple data sets. [^3] https://github.com/Garrafao/LSCDetection However, type-based models have more recently been outperformed by token-based contextualized embedding approaches (Kutuzov and Pivovarova, 2021; Zamora-Reina et al., 2022; Cassotti et al., 2023). Moreover, Duong et al. (2021) provide a repository to generate synthetic evaluation data for the related task of discourse shift detection with several model implementations not representing the state-of-the-art in LSCD. [^4] https://github.com/ruathudo/detangling-discourses
$G$
$G_{1}$
$G_{2}$
Figure 1: Word Usage Graph of English plane (left), subgraphs for first time period $G_{1}$ (middle) and for second time period $G_{2}$ (right). black/gray lines indicate high/low edge weights.
3 Tasks
LSCD can be seen as the combination of (at least) three lexical semantic tasks (Schlechtweg, 2023): (i) measurement of semantic proximity between word usages, (ii) clustering of the usages based on their semantic proximity, and (iii) estimation of semantic change labels from the obtained clusterings. Task (i) and (ii) corresponds to the lexicographic process of deriving word senses (Kilgarriff, 2007), while task (iii) measures LSC based on the derived word senses. The tasks need to be solved sequentially, in the order given above, as each is dependent on the output of the previous task, e.g., word usages can only be clustered once their semantic proximity has been estimated.
The three tasks are reflected in the human (e.g. Schlechtweg et al., 2020, 2021; Kutuzov et al., 2022) as well as the computational process (e.g. Giulianelli et al., 2020; Montariol et al., 2021; Laicher et al., 2021; Homskiy and Arefyev, 2022) of measuring lexical semantic change. [^5] However, it is not always obvious as annotation and modeling procedures often try to simplify or skip steps of this process. The first task is known as a standalone task under the name of "Word-in-Context" (WiC) (Pilehvar and Camacho-Collados, 2019) while the second task is known as the task of "Word Sense Induction" (WSI) (Schütze, 1998). A number of recently created LSCD data sets reflect all of these tasks and thus allow for evaluation of WiC and WSI along with LSCD models (i.a. Schlechtweg et al., 2021; Kurtyigit et al., 2021; Kutuzov et al., 2022; Zamora-Reina et al., 2022; Chen et al., 2023). [^6] The bulk of these data sets is listed at: https://www.ims.uni-stuttgart.de/data/wugs.
3.1 Word-in-Context
The Word-in-Context task is to determine if two words occurring in two text fragments have the same or different meanings. Usually two usages of the same word probably in different grammatical forms are given. [^7] However, there are datasets with examples consisting of usages of two different words that are similar in one of their meanings (Huang et al., 2012; Armendariz et al., 2020; Baldissin et al., 2022), and in the cross-lingual setup these two words and the corresponding text fragments are in different languages (Martelli et al., 2021). For example: {example} Von Hassel replied that he had such faith in the plane that he had no hesitation about allowing his only son to become a Starfighter pilot. {example} This point, where the rays pass through the perspective plane, is called the seat of their representation.
The WiC task is often framed as a binary classification task. For instance, the WiC (Pilehvar and Camacho-Collados, 2019) and MCL-WiC (Martelli et al., 2021) datasets contain binary labels and employ accuracy as the main evaluation metric. Alternatively, USim (Erk et al., 2013), SCWS (Huang et al., 2012) and CoSimLex (Armendariz et al., 2020) were labeled with non-binary similarity scores and promote a graded formulation of the task. In this formulation, a WiC model shall produce scores that are similar to the human scores, or at least rank the pairs of usages similarly. Spearman’s and Pearson’s correlation coefficients are employed as evaluation metrics in this case.
Following the DURel annotation framework (Schlechtweg et al., 2024), during the annotation process of most LSCD datasets human annotators were essentially solving the graded WiC task, i.e. they annotated the similarity of two usages of the same word on a scale. This provides data for evaluation of WiC models that can serve as a part of LSCD models. In diachronic LSCD datasets there are pairs of word usages extracted from two documents belonging to distant time periods making usages in these pairs very different orthographically, grammatically, and thematically, even when the target word has the same meaning. This might be challenging for models trained on traditional WiC datasets, which often contain examples from the same time period. Our benchmark helps to analyze how sensitive WiC models are to this shift in time period by comparing their performance on pairs of usages extracted from the old, the new or both corpora.
3.2 Word Sense Induction
The Word Sense Induction task is to infer which senses a given target word has based only on its usages in an unlabeled corpus. It is usually framed as a clustering task where a model shall cluster a given set of usages of the same target word probably in different grammatical forms into clusters corresponding to the senses of this word. Unlike the more popular Word Sense Disambiguation task, in WSI no sense inventory is given to the model and the number of senses of the target word is not known as well. The most widespread formulation of WSI assumes that each word usage has one and only one sense, thus, requires hard clustering, i.e. assigning each usage to a single cluster (e.g. SemEval 2010 Task 14 (Manandhar and Klapaftis, 2009) among many others). An alternative is modeling word meaning in context as a mixture of (not mutually exclusive) sense labels (e.g. SemEval 2013 Task 13 (Jurgens and Klapaftis, 2013)). The latter dataset contains examples with several senses assigned to a single word usage, thus, requiring soft clustering approaches.
Data set LGS $\mathbf{n}$ N/V/A $\mathbf{|U|}$ AN JUD Task $\mathbf{t_{1}}$ $\mathbf{t_{2}}$ Reference Version DWUG DE 48 32/14/2 178 8 37k WiC, WSI, LSCD (B,G,C) 1800–1899 1946–1990 Schlechtweg et al. (2021) 2.2.0 DWUG EN 40 36/4/0 189 9 29k WiC, WSI, LSCD (B,G,C) 1810–1860 1960–2010 Schlechtweg et al. (2021) 2.0.1 DWUG SV 40 31/6/3 168 5 20k WiC, WSI, LSCD (B,G,C) 1790–1830 1895–1903 Schlechtweg et al. (2021) 2.0.1 DWUG ES 100 51/24/25 40 12 62k WiC, WSI, LSCD (B,G,C) 1810–1906 1994–2020 Zamora-Reina et al. (2022) 4.0.0 DiscoWUG DE 75 39/16/20 49 8 24k WiC, WSI, LSCD (B,G,C) 1800–1899 1946–1990 Kurtyigit et al. (2021) 1.1.1 RefWUG DE 22 15/1/6 19 5 4k WiC, WSI, LSCD (B,G,C) 1750–1800 1850–1900 Schlechtweg (2023) 1.1.0 NorDiaChange1 NO 40 40/0/0 21 3 14k WiC, WSI, LSCD (B,G,C) 1929–1965 1970–2013 Kutuzov et al. (2022) 1.0.0 NorDiaChange2 NO 40 40/0/0 21 3 15k WiC, WSI, LSCD (B,G,C) 1980–1990 2012–2019 Kutuzov et al. (2022) 1.0.0 DURel DE 22 15/1/6 104 5 6k WiC, LSCD (C) 1750–1800 1850–1900 Schlechtweg et al. (2018) 3.0.0 SURel DE 22 19/3/0 104 4 5k WiC, LSCD (C) general domain Hätty et al. (2019) 3.0.0 RuSemShift1 RU 71 65/6/0 119 5 21k WiC, LSCD (C) 1682–1916 1918–1990 Rodina and Kutuzov (2020) 2.0.0 RuSemShift2 RU 69 57/12/0 105 5 18k WiC, LSCD (C) 1918–1990 1991–2016 Rodina and Kutuzov (2020) 2.0.0 RuShiftEval1 RU 111 111/0/0 60 3 10k WiC, LSCD (C) 1682–1916 1918–1990 Kutuzov and Pivovarova (2021) 2.0.0 RuShiftEval2 RU 111 111/0/0 60 3 10k WiC, LSCD (C) 1918–1990 1991–2016 Kutuzov and Pivovarova (2021) 2.0.0 RuShiftEval3 RU 111 111/0/0 60 3 10k WiC, LSCD (C) 1682–1916 1991–2016 Kutuzov and Pivovarova (2021) 2.0.0
Table 1: Overview data sets. LGS = language, $n$ = no. of target words, N/V/A = no. of nouns/verbs/adjectives, $|U|$ = avg. no. usages per word, AN = no. of annotators, JUD = total no. of judged usage pairs, Task = possible evaluation tasks, $t_{1}$, $t_{2}$ = time period 1/2, Reference = data set reference paper, Version = version used for experiments.
3.3 Lexical Semantic Change Detection
Lexical Semantic Change Detection is a general name for several tasks dealing with analysis of different properties of a word related to changes in its meaning over time. [^8] In this work we primarily discuss the diachronic setup studying the changes over time, but most discussions also generalize to the synchronic setup which studies how word meaning depends on text genre, topic or other factors rather than time period. In these tasks a list of target words are given and two time periods are specified, an old and a new one. [^9] Some datasets contain data for more than two time periods, but still the comparisons are made for each pair of time periods independently. Each time period is represented by an unlabeled corpus or a pre-selected set of usages for each target word.
The binary change task (Schlechtweg et al., 2020; Zamora-Reina et al., 2022) asks if the set of senses of a given word is the same for two time periods. It assumes that word meaning in a particular time period can be described as a set of discrete and mutually exclusive senses observed in the corresponding corpus. This task can be viewed as a task of binary classification of words. More specific versions of this task are the sense loss and the sense gain tasks asking if a word has lost any of its senses or obtained new senses between two time periods (Zamora-Reina et al., 2022).
The JSD task has the same assumptions about word meaning, but instead of binary classification it requires ranking a given list of words according to changes in their sense frequency distributions. The rank of a word is determined by the Jensen–Shannon Distance between two probability distributions over word senses $P(sense|w,t_{old})$ and $P(sense|w,t_{new})$, one for the older time period and another for the newer one (Schlechtweg et al., 2020; Zamora-Reina et al., 2022).
Finally, the COMPARE task requires ranking a given set of words according to the average similarity by meaning between old and new usages of each word (Schlechtweg et al., 2018; Kutuzov and Pivovarova, 2021; Zamora-Reina et al., 2022). The COMPARE task can be reduced to the graded WiC task for the pairs of usages consisting of one old and one new usage, the final word scores can be obtained by averaging WiC scores of these pairs.
All the presented LSCD tasks except for the COMPARE task require revealing which senses each target word has in each time period and comparing either two sets of senses corresponding to the old and the new time periods, or two frequency distributions over these senses. Thus, for solving these tasks it is reasonable to follow the lexicographic process first and employ WSI methods to cluster all usages of a word according to its senses, then perform some analysis of the obtained clusters.
Figure 2: Token-based LSCD pipelines and their evaluation.
4 Datasets
Table 1 shows all data sets currently integrated into the benchmark. All data sets have in common that they are based on human WiC judgments of word usage pairs (such as examples (3.1,3.1) from above) on the ordinal DURel scale from 1 to 4, where 1 means semantically unrelated and 4 means identical (Schlechtweg et al., 2018). They also share the use of diachronic data. [^10] With the exception of SURel comparing usages from different domains rather than time periods (Hätty et al., 2019).
The data sets then fall into two main categories: (i) data sets representing annotated judgments in a sparsely connected graph (Word Usage Graph, find an example in Figure 1), clustering these with a variation of correlation clustering (Bansal et al., 2004; Schlechtweg et al., 2021), and deriving LSC labels by comparing the two time-specific cluster frequency distributions (Schlechtweg et al., 2020). These data sets, displayed in the upper part of Table 1, apply the full lexicographic process and thus allow for full evaluation on all tasks mentioned in Section 3. (ii) Data sets skipping the clustering step and hence only allowing for evaluation on the WiC and COMPARE tasks. These are shown in the lower part of Table 1. Apart from the difference in tasks they support, the data sets have strongly varying properties in terms of language, number of target words, POS distribution, number of usages per target word and number of human judgments.
5 Evaluation Procedures
Figure 2 shows the structure of our benchmark. It summarizes token-based approaches to LSCD and shows how their components and whole pipelines can be evaluated using our benchmark.
The central part shows the WSI-based approach to LSCD. It relies on WSI methods which cluster word usages based either on their contextualized embeddings or pairwise similarities between them calculated by a WiC model. If a WiC model is involved, we can evaluate it separately on all datasets containing human-labeled pairs of word usages by feeding these pairs and comparing the model predictions with the human labels. Spearman’s and Person’s correlation coefficients that compare rankings or scores predicted by humans and the WiC model are employed as metrics for the WiC task.
A WSI method can also be evaluated as a whole by running it on a set of gold usages of each word, i.e. usages that have sense labels obtained either directly from human annotators, or by clustering word usage graphs. Clusterings obtained by the WSI method are compared against sense labels using Adjusted Rand Index (Hubert and Arabie, 1985) as a the main metric.
Finally, we can evaluate the whole LSCD pipelines using the standard LSCD metrics, i.e. F1-score for the binary classification tasks or Spearman’s correlation with the gold word ranking for the JSD and COMPARE tasks.
6 Conclusion
In this work we have presented a new benchmark for evaluation of token-based LSCD models. The procedures are implemented that can evaluate both whole LSCD solutions and their separate components that solve the WiC and WSI subtasks. A variety of LSCD datasets are integrated in the benchmark allowing thorough evaluation on 5 languages and diverse historical epochs.
Acknowledgments
Dominik Schlechtweg and Shafqat Mumtaz Virk have been funded by the research program ‘Change is Key!’ supported by Riksbankens Jubileumsfond (under reference number M21-0021). Nikolay Arefyev has received funding from the European Union’s Horizon Europe research and innovation program under Grant agreement No 101070350 (HPLT). Thanks to Andres Cabero and Kuan-Yu Lin for contributing code to the repository.