### v2

# The LSCD Benchmark: a Testbed for Diachronic Word Meaning Tasks

Abstract
Lexical Semantic Change Detection (LSCD) is a complex, lemma-level task, which is usually operationalized based on two subsequently applied usage-level tasks: First, Word-in-Context (WiC) labels are derived for pairs of usages. Then, these labels are represented in a graph on which Word Sense Induction (WSI) is applied to derive sense clusters. Finally, LSCD labels are derived by comparing sense clusters over time. This modularity is reflected in most LSCD datasets and models. It also leads to a large heterogeneity in modeling options and task definitions, which is exacerbated by a variety of dataset versions, preprocessing options and evaluation metrics. This heterogeneity makes it difficult to evaluate models under comparable conditions, to choose optimal model combinations or to reproduce results. Hence, we provide a benchmark repository standardizing LSCD evaluation. Through transparent implementation results become easily reproducible and by standardization different components can be freely combined. The repository reflects the task’s modularity by allowing model evaluation for WiC, WSI and LSCD. This allows for careful evaluation of increasingly complex model components providing new ways of model optimization. We use the implemented benchmark to conduct a number of experiments with recent models and systematically improve the state-of-the-art.
Keywords: Lexical Semantic Change Detection, Word-in-Context, Word Sense Induction, Word Usage Graphs, Diachronic, Word Meaning
[…]
The LSCD Benchmark: a Testbed for Diachronic Word Meaning Tasks
Dominik Schlechtweg1, Sachin Yadav1, Nikolay Arefyev2
1University of Stuttgart, 2University of Oslo
first.last@ims.uni-stuttgart.de, nikolare@uio.no
Abstract content
1.  Introduction
Lexical Semantic Change Detection (LSCD) is a field of NLP that studies methods automating the analysis of changes in word meanings over time. In recent years, this field has seen much development in terms of models, datasets and tasks (Schlechtweg, 2023). LSCD is a complex, lemma-level task, which is usually operationalized based on two subsequently applied usage-level tasks: First, Word-in-Context (WiC) labels are derived for pairs of usages. Then, these labels are represented in a graph on which Word Sense Induction (WSI) is applied to derive sense clusters. Finally, LSCD labels are derived by comparing sense clusters over time. This modularity is reflected in most LSCD datasets and models. It also leads to a large heterogeneity in modeling options and task definitions, which is exacerbated by a variety of dataset versions, preprocessing options and evaluation metrics. This heterogeneity makes it difficult to evaluate models under comparable conditions, to choose optimal model combinations or to reproduce results.
In order to handle this heterogeneity, we think that a shared testbed with a common evaluation setup is needed. Hence, we present a benchmark repository implementing evaluation procedures for models on most available LSCD datasets. [^1] Find the code at https://github.com/Garrafao/LSCDBenchmark. The benchmark exploits the modularity of the meta task LSCD by allowing for evaluation of the subtasks WiC and WSI on the same datasets. It can be assumed that performance on the subtasks directly determines performance on the meta task. We aim to stimulate transfer between the fields of WiC, WSI and LSCD by providing a repository allowing for evaluation on all these tasks with shared model components.
We hope that the resulting benchmark by standardizing the evaluation of LSCD models and providing models with SOTA performance can serve as a starting point for researchers to develop and improve models. The benchmark allows for a wide application and testing of models by focusing on multilingual models and their evaluation on several languages.
2.  Related Work
A number of recently created LSCD datasets apply WiC and WSI in the annotation process (cf. Section 3) and thus allow for evaluation of WiC and WSI along with LSCD models (i.a. Schlechtweg et al., 2021; Kurtyigit et al., 2021; Kutuzov et al., 2022; Zamora-Reina et al., 2022; Chen et al., 2023). [^2] The bulk of these datasets is listed at: https://www.ims.uni-stuttgart.de/data/wugs. There are also a number of datasets omitting WSI, but allowing for WiC and LSCD evaluation (i.a. Schlechtweg et al., 2018; Rodina and Kutuzov, 2020; Kutuzov and Pivovarova, 2021), or datasets omitting the WiC allowing for WSI and LSCD evaluation (i.a. Basile et al., 2020; Cook et al., 2014), or datasets allowing purely for WiC evaluation (Loureiro et al., 2022). Above that, there is a number of datasets allowing for synchronic evaluation of WiC (i.a. Pilehvar and Camacho-Collados, 2019; Trott and Bergen, 2021) or WSI (i.a. Langone et al., 2004; Hovy et al., 2006) or both (Erk et al., 2013; Aksenova et al., 2022), which can be exploited e.g. by simulating LSCD labels (Rosenfeld and Erk, 2018; Dubossarsky et al., 2019; Schlechtweg and Schulte im Walde, 2020). Note also that there is no restriction on the source or strategy for sampling word usages, i.e., they do not necessarily have to be sampled from different time periods, but could also be sampled from different text genres, domains, dialects or even languages. Hence, datasets reflecting meaning divergences between these text categories can be integrated into the benchmark as well (i.a. Hätty et al., 2019; Baldissin et al., 2022).
So far, there is no comprehensive LSCD benchmark, implementing state-of-the-art models on (human-annotated) high-quality evaluation data from multiple languages and multiple time periods. The leaderboards of several standalone shared tasks can be seen as small-scale benchmarks without common model implementation (Ahmad et al., 2020; Basile et al., 2020; Kutuzov and Pivovarova, 2021; Zamora-Reina et al., 2022; Fedorova et al., 2024) with the SemEval task being the most diverse with four languages (Schlechtweg et al., 2020). Schlechtweg et al. (2019) provide a comprehensive repository of type-based modeling approaches to LSCD with evaluation pipelines on multiple datasets. [^3] https://github.com/Garrafao/LSCDetection However, type-based models have more recently been outperformed by token-based contextualized embedding approaches (Kutuzov and Pivovarova, 2021; Zamora-Reina et al., 2022; Cassotti et al., 2023). Periti and Tahmasebi (2024) perform a systematic comparison, but do not provide a flexible model implementation and do not evaluate on some of the most recent SOTA models or datasets. Moreover, Duong et al. (2021) provide a repository to generate synthetic evaluation data for the related task of discourse shift detection with several model implementations not representing the state-of-the-art in LSCD. [^4] https://github.com/ruathudo/detangling-discourses
$G$
$G_{1}$
$G_{2}$
Figure 1: Word Usage Graph of English plane (left), subgraphs for first time period $G_{1}$ (middle) and for second time period $G_{2}$ (right). black/gray lines indicate high/low edge weights.
3.  Tasks
LSCD can be seen as the combination of (at least) three lexical semantic tasks (Schlechtweg, 2023): (i) measurement of semantic proximity between word usages, (ii) clustering of the usages based on their semantic proximity, and (iii) estimation of semantic change labels from the obtained clusterings. Task (i) and (ii) corresponds to the lexicographic process of deriving word senses (Kilgarriff, 2007), while task (iii) measures LSC based on the derived word senses. The tasks need to be solved sequentially, in the order given above, as each is dependent on the output of the previous task, e.g., word usages can only be clustered once their semantic proximity has been estimated.
The three tasks are reflected in the human (e.g. Schlechtweg et al., 2020, 2021; Kutuzov et al., 2022) as well as the computational process (e.g. Giulianelli et al., 2020; Montariol et al., 2021; Laicher et al., 2021; Homskiy and Arefyev, 2022) of measuring lexical semantic change. [^5] However, it is not always obvious as annotation and modeling procedures often try to simplify or skip steps of this process. The first task is known as a standalone task under the name of "Word-in-Context" (WiC) (Pilehvar and Camacho-Collados, 2019) while the second task is known as the task of "Word Sense Induction" (WSI) (Schütze, 1998). A number of recently created LSCD datasets reflect all of these tasks and thus allow for evaluation of WiC and WSI along with LSCD models (i.a. Schlechtweg et al., 2021; Kurtyigit et al., 2021; Kutuzov et al., 2022; Zamora-Reina et al., 2022; Chen et al., 2023). [^6] The bulk of these datasets is listed at: https://www.ims.uni-stuttgart.de/data/wugs.
3.1.  Word-in-Context
The Word-in-Context task is to determine if two words occurring in two text fragments have the same or different meanings. Usually two usages of the same word probably in different grammatical forms are given. [^7] However, there are datasets with examples consisting of usages of two different words that are similar in one of their meanings (Huang et al., 2012; Armendariz et al., 2020; Baldissin et al., 2022), and in the cross-lingual setup these two words and the corresponding text fragments are in different languages (Martelli et al., 2021). For example: {example} Von Hassel replied that he had such faith in the plane that he had no hesitation about allowing his only son to become a Starfighter pilot. {example} This point, where the rays pass through the perspective plane, is called the seat of their representation. The WiC task is often framed as a binary classification task. For instance, the WiC (Pilehvar and Camacho-Collados, 2019) and MCL-WiC (Martelli et al., 2021) datasets contain binary labels and employ accuracy as the main evaluation metric. Alternatively, USim (Erk et al., 2013), SCWS (Huang et al., 2012) and CoSimLex (Armendariz et al., 2020) were labeled with non-binary semantic proximity scores and promote a graded formulation of the task. In this formulation, a WiC model shall produce scores that are similar to the human scores, or at least rank the pairs of usages similarly. Spearman’s and Pearson’s correlation coefficients are employed as evaluation metrics in this case. More recently, Schlechtweg et al. (2025) provided an ordinal formulation of the graded WiC task evaluating with Krippendorff’s $\alpha$ (Krippendorff, 2018).
Following the DURel annotation framework (Schlechtweg et al., 2024b), during the annotation process of most LSCD datasets human annotators were essentially solving the graded WiC task, i.e., they annotated the semantic proximity of two usages of the same word on a scale. This provides data for evaluation of WiC models that can serve as a part of LSCD models. In diachronic LSCD datasets, there are pairs of word usages extracted from two documents belonging to distant time periods making usages in these pairs very different orthographically, grammatically, and thematically, even when the target word has the same meaning. This might be challenging for models trained on traditional WiC datasets, which often contain examples from the same time period. Our benchmark helps to analyze how sensitive WiC models are to this shift in time period by comparing their performance on pairs of usages extracted from the old, the new or both corpora.
3.2.  Word Sense Induction
The Word Sense Induction task is to infer which senses a given target word has based only on its usages in an unlabeled corpus. It is usually framed as a clustering task where a model shall cluster a given set of usages of the same target word probably in different grammatical forms into clusters corresponding to the senses of this word. Unlike the more popular Word Sense Disambiguation task, in WSI no sense inventory is given to the model and the number of senses of the target word is not known as well. The most widespread formulation of WSI assumes that each word usage has one and only one sense, thus, requires hard clustering, i.e. assigning each usage to a single cluster (e.g. SemEval 2010 Task 14 (Manandhar and Klapaftis, 2009) among many others). An alternative is modeling word meaning in context as a mixture of (not mutually exclusive) sense labels (e.g. SemEval 2013 Task 13 (Jurgens and Klapaftis, 2013)). The latter dataset contains examples with several senses assigned to a single word usage, thus, requiring soft clustering approaches.
Data set LGS $\mathbf{n}$ N/V/A $\mathbf{|U|}$ AN JUD Task $\mathbf{t_{1}}$ $\mathbf{t_{2}}$ Reference Version DWUG DE 50 32/14/2 178 8 63k WiC, WSI, LSCD (B,G,C) 1800–1899 1946–1990 Schlechtweg et al. (2021) 3.0.0 DWUG Res. DE 15 10/4/1 50 3 10k WiC, WSI, LSCD (B,G,C) 1800–1899 1946–1990 Schlechtweg et al. (2024a) 1.0.0 DWUG EN 46 40/6/0 191 13 29k WiC, WSI, LSCD (B,G,C) 1810–1860 1960–2010 Schlechtweg et al. (2021) 3.0.0 DWUG Res. EN 15 14/1/0 50 3 7K WiC, WSI, LSCD (B,G,C) 1810–1860 1960–2010 Schlechtweg et al. (2024a) 1.0.0 DWUG SV 44 32/5/7 171 13 20k WiC, WSI, LSCD (B,G,C) 1790–1830 1895–1903 Schlechtweg et al. (2021) 3.0.0 DWUG Res. SV 15 10/3/2 50 6 16k WiC, WSI, LSCD (B,G,C) 1790–1830 1895–1903 Schlechtweg et al. (2024a) 1.0.0 DWUG ES 100 51/24/25 40 12 62k WiC, WSI, LSCD (B,G,C) 1810–1906 1994–2020 Zamora-Reina et al. (2022) 4.0.2 DiscoWUG DE 75 39/16/20 49 8 24k WiC, WSI, LSCD (B,G,C) 1800–1899 1946–1990 Kurtyigit et al. (2021) 2.0.0 RefWUG DE 22 15/1/6 19 5 4k WiC, WSI, LSCD (B,G,C) 1750–1800 1850–1900 Schlechtweg (2023) 1.1.0 NorDiaChange1 NO 40 40/0/0 21 3 14k WiC, WSI, LSCD (B,G,C) 1929–1965 1970–2013 Kutuzov et al. (2022) 1.0.0 NorDiaChange2 NO 40 40/0/0 21 3 15k WiC, WSI, LSCD (B,G,C) 1980–1990 2012–2019 Kutuzov et al. (2022) 1.0.0 ChiWUG ZH 40 10/22/8 40 4 61k WiC, WSI, LSCD (B,G,C) 954-1978 1979-2003 Chen et al. (2023) 1.0.0 DWUG IT 5k WiC, WSI, LSCD (B,G,C) 1948-1970 1990-2014 Cassotti et al. (2024) 3.0.0 DURel DE 22 15/1/6 104 5 6k WiC, LSCD (C) 1750–1800 1850–1900 Schlechtweg et al. (2018) 3.0.0 SURel DE 22 19/3/0 104 4 5k WiC, LSCD (C) general domain Hätty et al. (2019) 3.0.0 RuSemShift1 RU 71 65/6/0 119 5 21k WiC, LSCD (C) 1682–1916 1918–1990 Rodina and Kutuzov (2020) 2.0.0 RuSemShift2 RU 69 57/12/0 105 5 18k WiC, LSCD (C) 1918–1990 1991–2016 Rodina and Kutuzov (2020) 2.0.0 RuShiftEval1 RU 111 111/0/0 60 3 10k WiC, LSCD (C) 1682–1916 1918–1990 Kutuzov and Pivovarova (2021) 2.0.0 RuShiftEval2 RU 111 111/0/0 60 3 10k WiC, LSCD (C) 1918–1990 1991–2016 Kutuzov and Pivovarova (2021) 2.0.0 RuShiftEval3 RU 111 111/0/0 60 3 10k WiC, LSCD (C) 1682–1916 1991–2016 Kutuzov and Pivovarova (2021) 2.0.0
Table 1: Overview datasets. LGS = language, $n$ = no. of target words, N/V/A = no. of nouns/verbs/adjectives, $|U|$ = avg. no. usages per word, AN = no. of annotators, JUD = total no. of judged usage pairs, Task = possible evaluation tasks, $t_{1}$, $t_{2}$ = time period 1/2, Reference = data set reference paper, Version = version used for experiments.
$3.$  Lexical Semantic Change Detection
Lexical Semantic Change Detection is a general name for several tasks dealing with analysis of different properties of a word related to changes in its meaning over time. [^8] In this work we primarily discuss the diachronic setup studying the changes over time, but most discussions also generalize to the synchronic setup which studies how word meaning depends on text genre, topic or other factors rather than time period. In these tasks a list of target words are given and two time periods are specified, an old and a new one. [^9] Some datasets contain data for more than two time periods, but still the comparisons are made for each pair of time periods independently. Each time period is represented by an unlabeled corpus or a pre-selected set of usages for each target word.
The binary change task (Schlechtweg et al., 2020; Zamora-Reina et al., 2022) asks if the set of senses of a given word is the same for two time periods. It assumes that word meaning in a particular time period can be described as a set of discrete and mutually exclusive senses observed in the corresponding corpus. This task can be viewed as a task of binary classification of words. More specific versions of this task are the sense loss and the sense gain tasks asking if a word has lost any of its senses or obtained new senses between two time periods (Zamora-Reina et al., 2022).
The JSD (or Graded Change) task has the same assumptions about word meaning, but instead of binary classification it requires ranking a given list of words according to changes in their sense frequency distributions. The rank of a word is determined by the Jensen–Shannon Distance between two probability distributions over word senses $P(sense|w,t_{old})$ and $P(sense|w,t_{new})$, one for the older time period and another for the newer one (Schlechtweg et al., 2020; Zamora-Reina et al., 2022).
Finally, the COMPARE task requires ranking a given set of words according to the average proximity by meaning between old and new usages of each word (Schlechtweg et al., 2018; Kutuzov and Pivovarova, 2021; Zamora-Reina et al., 2022). The COMPARE task can be reduced to the graded WiC task for the pairs of usages consisting of one old and one new usage, the final word scores can be obtained by averaging WiC scores of these pairs.
All the presented LSCD tasks except for the COMPARE task require revealing which senses each target word has in each time period and comparing either two sets of senses corresponding to the old and the new time periods, or two frequency distributions over these senses. Thus, for solving these tasks it is reasonable to follow the lexicographic process first and employ WSI methods to cluster all usages of a word according to its senses, then perform some analysis of the obtained clusters.
Figure 2: Token-based LSCD pipelines and their evaluation.
4.  Datasets
Table 1 shows all datasets currently integrated into the benchmark. All datasets have in common that they are based on human WiC judgments of word usage pairs (such as examples (3.1,3.1) from above) on the ordinal DURel scale from 1 to 4, where 1 means semantically unrelated and 4 means identical (Schlechtweg et al., 2018). They also share the use of diachronic data. [^10] With the exception of SURel comparing usages from different domains rather than time periods (Hätty et al., 2019).
The datasets then fall into two main categories: (i) datasets representing annotated judgments in a sparsely connected graph (Word Usage Graph, find an example in Figure 1), clustering these with a variation of correlation clustering (Bansal et al., 2004; Schlechtweg et al., 2021), and deriving LSC labels by comparing the two time-specific cluster frequency distributions (Schlechtweg et al., 2020). These datasets, displayed in the upper part of Table 1, apply the full lexicographic process and thus allow for full evaluation on all tasks mentioned in Section 3. (ii) datasets skipping the clustering step and hence only allowing for evaluation on the WiC and COMPARE tasks. These are shown in the lower part of Table 1. Apart from the difference in tasks they support, the datasets have strongly varying properties in terms of language, number of target words, POS distribution, number of usages per target word and number of human judgments.
5.  Evaluation Procedures
Figure 2 shows the structure of our benchmark. It summarizes token-based approaches to LSCD and shows how their components and whole pipelines can be evaluated using our benchmark.
The central part shows the WSI-based approach to LSCD. It relies on WSI methods which cluster word usages based either on their contextualized embeddings or pairwise similarities between them calculated by a WiC model. If a WiC model is involved, we can evaluate it separately on all datasets containing human-labeled pairs of word usages by feeding these pairs and comparing the model predictions with the human labels. Spearman’s and Pearson’s correlation coefficients that compare rankings or scores predicted by humans and the WiC model are employed as metrics for the WiC task.
A WSI method can also be evaluated as a whole by running it on a set of gold usages of each word, i.e. usages that have sense labels obtained either directly from human annotators, or by clustering word usage graphs. Clusterings obtained by the WSI method are compared against sense labels using Adjusted Rand Index (Hubert and Arabie, 1985) as a the main metric.
Finally, we can evaluate the whole LSCD pipelines using the standard LSCD metrics, i.e. F1-score for the binary classification tasks or Spearman’s correlation with the gold word ranking for the JSD and COMPARE tasks.
We introduce standard splits for each dataset on the lemma level, i.e., certain target words are assigned to train/dev/test. We further provide possibility to evaluate on the previously introduced standard split from the CoMeDi shared task (Schlechtweg et al., 2025).
6.  Models
The most straightforward modeling approach for LSCD models is to follow the 3-level annotation approach of WUGs described in Section 3. Given a target word, a basic model retrieves uses of this word from an old and a new corpus, then clusters them in order to infer word senses, and finally analyzes the obtained clusters to make conclusions about changes in word senses between two time periods. This approach is appealing because if word senses are inferred correctly, then an exact description of how word senses changed, as well as the exact predictions for all LSCD tasks are easy to obtain. Also the procedure basically automates the annotation procedure of various LSCD datasets, which is a reasonable way of getting predictions that correspond to the ground truth well. The benchmark implementation as depicted in Figure 2 thus follows this basic structure. The inputs and the outputs of each component are specified in round brackets, while the hyperparameters are underlined in the corresponding descriptions. The implemented components on each of these levels will be described below. It is important to note that current SOTA models for graded change (Giulianelli et al., 2020; Arefyev et al., 2021a; Cassotti et al., 2023) skip the clustering step and model graded change directly from Word-in-Context predictions of proximity between word usages or their underlying vectors by aggregating these time-wise (see below).
Figure 3: Result overview on Graded Change (left) and COMPARE score (right). XLD = XL-DURel, XLM = XL-LEXEME, MCL/MCLen = DeepMistake checkpoints, TH = thresholded, APD = Average Pairwise Distance, DIA = DiaSense, COS = Cosine distance between average embeddings.
Retrieve uses
(word, corpus → uses). Given a corpus and a target word, this component retrieves all uses of the target word in all of its grammatical forms from the given corpus. Optionally, N uses may be sampled if there are more than that in order to reduce the following computations.
For intrinsic evaluation, instead of retrieving uses, golden uses may be taken, i.e., the uses which were shown to the annotators during dataset construction. This eliminates possible disagreements due to some rare senses of the target word sampled during the dataset annotation procedure but not sampled by the model during evaluation, or vice versa.
Build contextualized embeddings
(uses → embeddings). We employ BERT or XLM-R masked language model to obtain the contextualized representations of the given target word in the given text fragment (Devlin et al., 2019; Conneau et al., 2019). [^11] In principle, the benchmark supports all model checkpoints from huggingface. The simplest and most popular option is taking the outputs of the last Transformer layer (before the MLM head) on the positions of the target word and average those outputs (mean pooling). We can apply other subword poolings, max pooling or first pooling (take the outputs from the first subword). A more general implementation combines the outputs from several layers. Again, the simplest option is just averaging them. The experiments with BERT without fine-tuning presented in (Devlin et al., 2019) for the NER task suggest that it may be better to concatenate the outputs of the last four layers instead of averaging them. Thus, the layer aggregation function is a hyperparameter and we select its value among averaging and concatenation. Additionally, the benchmark supports XL-LEXEME, a WiC-fine-tuned version of XLM-R (Cassotti et al., 2023) and the similar XL-DURel which is instead fine-tuned for ordinal WiC (Yadav and Schlechtweg, 2025).
Generate pairs
(uses → pairs of uses). Pairs of uses of the specified type are generated. For the COMPARE type, each pair contains a use from the old corpus and a use from the new corpus. This type is ideal for the COMPARE task in which one needs to estimate the average similarity between old and new examples. If type is ALL, then all possible pairs are generated. This is useful to provide more information for the following clustering step.
For intrinsic evaluation, the golden pairs, i.e., those pairs that were annotated by humans during the dataset construction process, may be directly fed to the following components.
WiC
(pairs of uses → pairwise scores). We consider two different approaches to the Word-in-Context task. The first approach builds the contextualized embeddings for all uses with the aforementioned component, it inherits all the corresponding hyperparameters. Then it employs one of the distance functions to calculate distances between the contextualized embeddings of two uses in each pair. The euclidean, manhattan and cosine distances are currently supported.
The second approach employs a binary classifier implemented as a neural network that jointly processes a pair of word usages and returns the probability that the meaning is the same. We treat this probability as the similarity between word occurrences. We experimented with several such models that were part of the SOTA or near-SOTA solutions proposed by the DeepMistake team in the RuShiftEval and LSCDiscovery shared tasks (Arefyev et al., 2021a; Homskiy and Arefyev, 2022) or showed good performance on recent ordinal WiC shared task data (Yadav and Schlechtweg, 2025; Schlechtweg et al., 2025).
We further provide the possibility of discretizing graded WiC predictions at specified thresholds. By default, the thresholds from Yadav and Schlechtweg (2025) are used.
Clustering
(pairwise scores → clusters or contextualized embeddings → clusters). The goal of this step is to discover all senses of the target word occurring in the old, or the new corpus, or both of them, i.e., solve the WSI task for all uses retrieved from both corpora. Those clustering algorithms that can accept a matrix of similarities or distances between objects instead of object vectors can be applied to both types of inputs to the clustering component (pairwise scores, embeddings). However, some clustering algorithms require raw object vectors and can be applied to the contextualized embeddings, but not to the pairwise scores. For instance, K-Means calculates cluster centroids and needs object vectors to do that. The benchmark currently supports correlation clustering (Bansal et al., 2004) operating on pairwise scores. We chose this algorithm as most gold annotations were clustered with this approach (e.g. Schlechtweg et al., 2020).
Cluster measures
(clusters → LSCD predictions). To solve the binary LSCD tasks, after WSI we search for those clusters that contain only new or only old examples. Those clusters are viewed as novel or lost senses correspondingly, and the binary labels are predicted accordingly. Alternatively, we search for the clusters with at least M new examples and at most K old examples or vice versa in order to align with the annotation procedure of some LSCD datasets. To solve the JSD task, for old and new uses separately we estimate the probability distribution over senses of the target word as the proportions of its uses ended in each cluster. Then JSD between two probability distributions is calculated. Finally, to solve the COMPARE task we average the indicators that two uses ended in the same cluster for all pairs of uses of the COMPARE type.
Figure 4: Result overview on WiC.
Aggregate measures
(pairwise scores → LSCD predictions or contextualized embeddings → LSCD predictions). These skip the clustering step by aggregating WiC predictions directly. The most successful of these is Average Pairwise Distance (APD), which simply averages the distances or similarities between word usages from different time periods (COMPARE pairs) returned by some WiC model (Kutuzov and Giulianelli, 2020). This follows the calculation of the gold COMPARE scores, which is the average of human pairwise annotations. APD is sensitive to the polysemy or variation of the target word, which can lead to wrong predictions. For this, measures normalizing APD by a polysemy term were proposed, such as DiaSense (Beck, 2020). COS, in contrast, averages the contextualized embeddings of all old and all new uses separately producing two aggregated embeddings of the target word for each of two time periods, then calculate the distance between those two embeddings. It was previously under the names of PRT (Kutuzov and Giulianelli, 2020) and COS (Laicher et al., 2021). This distance is used as the prediction for both graded tasks.
7.  Experiments
We now use the benchmark the perform a number of experiments. We focus on Graded Change and COMPARE detection as these are the most widely approached LSCD tasks (Kutuzov and Pivovarova, 2021; Periti and Tahmasebi, 2024). Note that not all datasets provide both evaluation scores (see Table 1). All experiments are performed on the CoMeDi test split described in Section 1. We cannot test all models and configurations, thus we focus on SOTA models using aggregate change measures and design experiments to answer open research questions.
Which WiC model and which aggregate measure gives SOTA performance?
According to recent studies (Zamora-Reina et al., 2022; Periti and Tahmasebi, 2024; Zamora-Reina et al., 2025), DeepMistake and XL-LEXEME combined with APD compete for the SOTA on Graded Change and COMPARE detection. Previous studies have not directly compared these models though, or only on very limited data (Zamora-Reina et al., 2025). Hence, we perform a direct comparison guaranteeing a fair evaluation setup. We also include the recently published XL-DURel (Yadav and Schlechtweg, 2025), which has been shown to improve upon both models on ordinal WiC. In addition to the canonical datasets, we include a number of recently published datasets which have never be used for model evaluation: DWUG DE/EN/SV V3.0.0, DWUG DE/EN/SV resampled and DWUG IT. This is the most thorough model comparison done so far in the field of LSCD.
Figure 5: Result overview on dataset versions for bi-encoder models.
Figure 3 shows the results across the three SOTA WiC models, each combined with the common aggregate change measures APD, COS and DiaSense (see Section 6). Both, for Graded Change and COMPARE, clearly APD dominates which confirms previous results. However, there are slight advantages for the recently published XL-DURel model suggesting a new SOTA.
Does WiC prediction discretization improve results?
All above-reported measures based on aggregation of graded WiC predictions (either cosine similarity or same-sense probability). However, all WUG datasets are annotated on an ordinal (discrete) scale (see Section 4). Exactly predicting these ordinal values can be done by thresholding the graded predictions (Choppa et al., 2025). We hypothesize that discretizing WiC predictions to resemble human annotations helps for LSCD as ordinal judgments were used for ground-truth construction. Hence, in Figure 3 we report two models with APD and thresholding (XLD/XLM-TH-APD). Threshold parameters were taken from Yadav and Schlechtweg (2025). As we see, for XL-DURel, thresholding slightly helps, further pushing the SOTA, while for XL-LEXEME it has no effect or slightly hurts.
Which model gives SOTA on diachronic WiC? Does WiC determine LSCD?
A recent shared task has compared DeepMistake and XL-LEXEME for ordinal WiC showing a slight advantage for DeepMistake (Schlechtweg et al., 2025). [^12] Note that these models do not exactly correspond to the checkpoints we used. XL-DURel has further improved upon both models. Performance on the WiC task is a strongly influential factor for LSCD (Arefyev et al., 2021b). We compare model performance on these two levels to gain an understanding how strongly WiC determines LSCD performance. Figure 4 compares WiC models on the ordinal WiC task. The MCL checkpoint of DeepMistake does dominate XL-LEXEME confirming previous results, but both are outperformed by XL-DURel. Now compare this to the APD columns in Figure 3. Overall, WiC performance determines LSCD performance. Consider e.g. the dominance of XL-DURel on DWUG EN with graded change. However, there are notable exceptions such as DWUG SV where XL-DURel dominates on WiC but not LSCD. This shows that purely improving WiC ranking does not guarantee better LSCD performance. Possibly, the score distribution plays an important role when averaging values for aggregate measures.
Are model performances reproducible with more reliable data? What is the performance development on incrementally annotated datasets?
Many datasets have been annotated in incremental rounds of annotation (Schlechtweg et al., 2021, 2024a). Later rounds are supposed to yield higher data quality as graphs are more richly annotated. Schlechtweg et al. (2024a) suggest that previous model comparisons done on older dataset versions/less rounds should be repeated with the more reliable data (last round). We are the first to investigate model performance with the latest dataset versions. This will also allow us to investigate the impact of annotation rounds on performance. Figure 5 shows the performance of a selection of models on consecutive versions. For the top models, we can see that performance tends to increase with later versions, which is expected as quality should increase. However, there are cases where model performance strongly drops for newer versions, such as XLM-APD on DWUG EN. Further, the relative performance of models can strongly change depending on version, e.g. XLM-APD outperforms XLD-APD on DWUG EN for an older version while it is the other way around for the newer, more reliable version. This shows the risks of using unreliable ground-truth data and illustrates the need for the creation of benchmarks such as ours and continuous model reevaluation on additional and improved datasets.
Figure 6: Result selection on preprocessing. lemma = lemmatization, norm = normalization, raw = no preprocessing, token = tokenization, toklem = tokenization with target word substituted by lemma.
What is the impact of spelling variation on performance?
Laicher (2020) show that historical spelling variations can have a strong influence on BERT-based model performance. However, it is not clear how much this influences SOTA models. Hence, we test the influence of spelling normalization and lemmatization on two German datasets containing historical spelling variants in Figure 6. We can see that applying no preprocessing (raw) gives top performance for all models, with one exception (MCLen with toklem). This suggests that current base embedders are quite robust against spelling variation and do not need additional preprocessing. [^13] Note that ‘raw’ on these datasets already incorporates minimal preprocessing substituting a small number of historical characters with modern ones.
8.  Conclusion
In this work, we have presented a new benchmark for evaluation of token-based LSCD models. The procedures are implemented that can evaluate both whole LSCD solutions and their separate components that solve the WiC and WSI subtasks. A variety of LSCD datasets are integrated in the benchmark allowing thorough evaluation on various languages and diverse historical epochs. We used the benchmark to perform a number of experiments with recent models setting a new state-of-the-art in LSCD and providing a better general understanding on LSCD model evaluation and improvement. We hope that our benchmark will inspire further research in this field.
Limitations
In our evaluation, we did not evaluate any cluster-based models although these have a large potential for high performance Zamora-Reina et al. (2022); Schlechtweg et al. (2024c). However, the current state-of-the-art does not build on clustering. Hence, we focused our evaluation to give a more concise overview and leave the comparison to cluster-based models to future work.
Acknowledgments
Dominik Schlechtweg and Sachin Yadav have been funded by the research program ‘Change is Key!’ supported by Riksbankens Jubileumsfond (under reference number M21-0021). Nikolay Arefyev has received funding from the European Union’s Horizon Europe research and innovation program under Grant agreement No 101070350 (HPLT). Thanks to Andres Cabero, Kuan-Yu Lin and Arshan Seyed Dalili for contributing code to the repository. Thanks to Shafqat Mumtaz Virk for contributing to an earlier version of this paper.
9.  Bibliographical References
[…]
  • Ahmad et al. (2020) Adnan Ahmad, Kiflom Desta, Fabian Lang, and Dominik Schlechtweg. 2020. Shared task: Lexical semantic change detection in german.
  • Aksenova et al. (2022) Anna Aksenova, Ekaterina Gavrishina, Elisei Rykov, and Andrey Kutuzov. 2022. RuDSI: Graph-based word sense induction dataset for Russian. In Proceedings of TextGraphs-16: Graph-based Methods for Natural Language Processing, pages 77–88, Gyeongju, Republic of Korea. Association for Computational Linguistics.
  • Arefyev et al. (2021a) N. Arefyev, M. Fedoseev, V. Protasov, D. Homskiy, A. Davletov, and A. Panchenko. 2021a. Deepmistake: Which senses are hard to distinguish for a word-in-context model. In Computational linguistics and intellectual technologies, 20, page 16 – 30, Russian Federation.
  • Arefyev et al. (2021b) Nikolay Arefyev, Maksim Fedoseev, Vitaly Protasov, Daniil Homskiy, Adis Davletov, and Alexander Panchenko. 2021b. DeepMistake: Which senses are hard to distinguish for a word-in-context model. volume 2021-June, pages 16–30.
  • Armendariz et al. (2020) Carlos Santos Armendariz, Matthew Purver, Matej Ulčar, Senja Pollak, Nikola Ljubešić, and Mark Granroth-Wilding. 2020. CoSimLex: A resource for evaluating graded word similarity in context. In Proceedings of the Twelfth Language Resources and Evaluation Conference, pages 5878–5886, Marseille, France. European Language Resources Association.
  • Baldissin et al. (2022) Gioia Baldissin, Dominik Schlechtweg, and Sabine Schulte im Walde. 2022. DiaWUG: A Dataset for Diatopic Lexical Semantic Variation in Spanish. In Proceedings of the 13th Language Resources and Evaluation Conference, Marseille, France. European Language Resources Association.
  • Bansal et al. (2004) Nikhil Bansal, Avrim Blum, and Shuchi Chawla. 2004. Correlation clustering. Machine Learning, 56(1-3):89–113.
  • Basile et al. (2020) Pierpaolo Basile, Annalina Caputo, Tommaso Caselli, Pierluigi Cassotti, and Rossella Varvara. 2020. Overview of the EVALITA 2020 Diachronic Lexical Semantics (DIACR-Ita) Task. In Proceedings of the 7th evaluation campaign of Natural Language Processing and Speech tools for Italian (EVALITA 2020), Online. CEUR.org.
  • Beck (2020) Christin Beck. 2020. DiaSense at SemEval-2020 Task 1: Modeling sense change via pre-trained BERT embeddings. In Proceedings of the 14th International Workshop on Semantic Evaluation, Barcelona, Spain. Association for Computational Linguistics.
  • Cassotti et al. (2024) Pierluigi Cassotti, Pierpaolo Basile, and Nina Tahmasebi. 2024. DWUGs-IT: Extending and standardizing lexical semantic change detection for Italian. In Proceedings of the 10th Italian Conference on Computational Linguistics, Pisa, Italy, December 4 - December 6, 2024, CEUR Workshop Proceedings. CEUR-WS.org.
  • Cassotti et al. (2023) Pierluigi Cassotti, Lucia Siciliani, Marco de Gemmis, Giovanni Semeraro, and Pierpaolo Basile. 2023. Xl-lexeme: Wic pretrained model for cross-lingual lexical semantic change. In Proceedings of the 61th Annual Meeting of the Association for Computational Linguistics, Online. Association for Computational Linguistics.
  • Chen et al. (2023) Jing Chen, Emmanuele Chersoni, Dominik Schlechtweg, Jelena Prokic, and Chu-Ren Huang. 2023. ChiWUG: A graph-based evaluation dataset for Chinese lexical semantic change detection. In Proceedings of the 4th International Workshop on Computational Approaches to Historical Language Change, Singapore. Association for Computational Linguistics.
  • Choppa et al. (2025) Tejaswi Choppa, Michael Roth, and Dominik Schlechtweg. 2025. Predicting median, disagreement and noise label in ordinal word-in-context data. In Proceedings of Context and Meaning: Navigating Disagreements in NLP Annotation, pages 65–77, Abu Dhabi, UAE. International Committee on Computational Linguistics.
  • Conneau et al. (2019) Alexis Conneau, Kartikay Khandelwal, Naman Goyal, Vishrav Chaudhary, Guillaume Wenzek, Francisco Guzmán, Edouard Grave, Myle Ott, Luke Zettlemoyer, and Veselin Stoyanov. 2019. Unsupervised cross-lingual representation learning at scale. arXiv preprint arXiv:1911.02116.
  • Cook et al. (2014) Paul Cook, Jey Han Lau, Diana McCarthy, and Timothy Baldwin. 2014. Novel word-sense identification. In COLING, pages 1624–1635. ACL.
  • Devlin et al. (2019) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019. BERT: Pre-training of deep bidirectional transformers for language understanding. In Proceedings of the 2019 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long and Short Papers), pages 4171–4186, Minneapolis, Minnesota. Association for Computational Linguistics.
  • Dubossarsky et al. (2019) Haim Dubossarsky, Simon Hengchen, Nina Tahmasebi, and Dominik Schlechtweg. 2019. Time-Out: Temporal Referencing for Robust Modeling of Lexical Semantic Change. In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, pages 457–470, Florence, Italy. Association for Computational Linguistics.
  • Duong et al. (2021) Quan Duong, Lidia Pivovarova, and Elaine Zosa. 2021. Benchmarks for unsupervised discourse change detection. In Proceedings of the 6th International Workshop on Computational History (HistoInformatics 2021) co-located with ACM/IEEE Joint Conference on Digital Libraries 2021 (JCDL 2021), Online event, September 30-October 1, 2021, volume 2981 of CEUR Workshop Proceedings. CEUR-WS.org.
  • Erk et al. (2013) Katrin Erk, Diana McCarthy, and Nicholas Gaylord. 2013. Measuring word meaning in context. Computational Linguistics, 39(3):511–554.
  • Fedorova et al. (2024) Mariia Fedorova, Timothee Mickus, Niko Partanen, Janine Siewert, Elena Spaziani, and Andrey Kutuzov. 2024. AXOLOTL’24 shared task on multilingual explainable semantic change modeling. In Proceedings of the 5th Workshop on Computational Approaches to Historical Language Change, pages 72–91, Bangkok, Thailand. Association for Computational Linguistics.
  • Giulianelli et al. (2020) Mario Giulianelli, Marco del Tredici, and Raquel Fernández. 2020. Analysing lexical semantic change with contextualised word representations. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 3960–3973, Online. Association for Computational Linguistics.
  • Hätty et al. (2019) Anna Hätty, Dominik Schlechtweg, and Sabine Schulte im Walde. 2019. SURel: A gold standard for incorporating meaning shifts into term extraction. In Proceedings of the 8th Joint Conference on Lexical and Computational Semantics, pages 1–8, Minneapolis, MN, USA.
  • Homskiy and Arefyev (2022) Daniil Homskiy and Nikolay Arefyev. 2022. DeepMistake at LSCDiscovery: Can a multilingual word-in-context model replace human annotators? In Proceedings of the 3rd Workshop on Computational Approaches to Historical Language Change, pages 173–179, Dublin, Ireland. Association for Computational Linguistics.
  • Hovy et al. (2006) Eduard Hovy, Mitchell Marcus, Martha Palmer, Lance Ramshaw, and Ralph Weischedel. 2006. Ontonotes: The 90% solution. In Proceedings of the Human Language Technology Conference of the NAACL, Companion Volume: Short Papers, NAACL-Short ’06, pages 57––60, USA. Association for Computational Linguistics.
  • Huang et al. (2012) Eric Huang, Richard Socher, Christopher Manning, and Andrew Ng. 2012. Improving word representations via global context and multiple word prototypes. In Proceedings of the 50th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 873–882, Jeju Island, Korea. Association for Computational Linguistics.
  • Hubert and Arabie (1985) Lawrence Hubert and Phipps Arabie. 1985. Comparing partitions. Journal of Classification, 2(1):193–218.
  • Jurgens and Klapaftis (2013) David Jurgens and Ioannis Klapaftis. 2013. SemEval-2013 task 13: Word sense induction for graded and non-graded senses. In Second Joint Conference on Lexical and Computational Semantics (*SEM), Volume 2: Proceedings of the Seventh International Workshop on Semantic Evaluation (SemEval 2013), pages 290–299, Atlanta, Georgia, USA. Association for Computational Linguistics.
  • Kilgarriff (2007) Adam Kilgarriff. 2007. Word Senses, chapter 2. Springer.
  • Krippendorff (2018) Klaus Krippendorff. 2018. Content Analysis: An Introduction to Its Methodology. SAGE Publications.
  • Kurtyigit et al. (2021) Sinan Kurtyigit, Maike Park, Dominik Schlechtweg, Jonas Kuhn, and Sabine Schulte im Walde. 2021. Lexical Semantic Change Discovery. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers), Online. Association for Computational Linguistics.
  • Kutuzov and Giulianelli (2020) Andrey Kutuzov and Mario Giulianelli. 2020. UiO-UvA at SemEval-2020 Task 1: Contextualised Embeddings for Lexical Semantic Change Detection. In Proceedings of the 14th International Workshop on Semantic Evaluation, Barcelona, Spain. Association for Computational Linguistics.
  • Kutuzov and Pivovarova (2021) Andrey Kutuzov and Lidia Pivovarova. 2021. Rushifteval: a shared task on semantic shift detection for russian. Komp’yuternaya Lingvistika i Intellektual’nye Tekhnologii: Dialog conference.
  • Kutuzov et al. (2022) Andrey Kutuzov, Samia Touileb, Petter Mæhlum, Tita Enstad, and Alexandra Wittemann. 2022. NorDiaChange: Diachronic semantic change dataset for Norwegian. In Proceedings of the Thirteenth Language Resources and Evaluation Conference, pages 2563–2572, Marseille, France. European Language Resources Association.
  • Laicher (2020) Severin Laicher. 2020. Historical word sense clustering with deep contextualized word embeddings.
  • Laicher et al. (2021) Severin Laicher, Sinan Kurtyigit, Dominik Schlechtweg, Jonas Kuhn, and Sabine Schulte im Walde. 2021. Explaining and Improving BERT Performance on Lexical Semantic Change Detection. In Proceedings of the 16th Conference of the European Chapter of the Association for Computational Linguistics: Student Research Workshop, pages 192–202, Online. Association for Computational Linguistics.
  • Langone et al. (2004) Helen Langone, Benjamin R. Haskell, and George A. Miller. 2004. Annotating wordnet. In Proceedings of the Workshop Frontiers in Corpus Annotation at HLT-NAACL, Boston, MA, USA.
  • Loureiro et al. (2022) Daniel Loureiro, Aminette D’Souza, Areej Nasser Muhajab, Isabella A. White, Gabriel Wong, Luis Espinosa-Anke, Leonardo Neves, Francesco Barbieri, and Jose Camacho-Collados. 2022. TempoWiC: An evaluation benchmark for detecting meaning shift in social media. In Proceedings of the 29th International Conference on Computational Linguistics, pages 3353–3359, Gyeongju, Republic of Korea. International Committee on Computational Linguistics.
  • Manandhar and Klapaftis (2009) Suresh Manandhar and Ioannis Klapaftis. 2009. SemEval-2010 task 14: Evaluation setting for word sense induction & disambiguation systems. In Proceedings of the Workshop on Semantic Evaluations: Recent Achievements and Future Directions (SEW-2009), pages 117–122, Boulder, Colorado. Association for Computational Linguistics.
  • Martelli et al. (2021) Federico Martelli, Najla Kalach, Gabriele Tola, and Roberto Navigli. 2021. SemEval-2021 task 2: Multilingual and cross-lingual word-in-context disambiguation (MCL-WiC). In Proceedings of the 15th International Workshop on Semantic Evaluation (SemEval-2021), pages 24–36, Online. Association for Computational Linguistics.
  • Montariol et al. (2021) Syrielle Montariol, Matej Martinc, and Lidia Pivovarova. 2021. Scalable and interpretable semantic change detection. In 2021 Annual Conference of the North American Chapter of the Association for Computational Linguistics.
  • Periti and Tahmasebi (2024) Francesco Periti and Nina Tahmasebi. 2024. A systematic comparison of contextualized word embeddings for lexical semantic change. In Proceedings of the 2024 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 1: Long Papers), pages 4262–4282. Association for Computational Linguistics.
  • Pilehvar and Camacho-Collados (2019) Mohammad Taher Pilehvar and Jose Camacho-Collados. 2019. WiC: the word-in-context dataset for evaluating context-sensitive meaning representations. In Proceedings of the 2019 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long and Short Papers), pages 1267–1273, Minneapolis, Minnesota. Association for Computational Linguistics.
  • Rodina and Kutuzov (2020) Julia Rodina and Andrey Kutuzov. 2020. RuSemShift: a dataset of historical lexical semantic change in Russian. In Proceedings of the 28th International Conference on Computational Linguistics, pages 1037–1047, Barcelona, Spain (Online). International Committee on Computational Linguistics.
  • Rosenfeld and Erk (2018) Alex Rosenfeld and Katrin Erk. 2018. Deep neural models of semantic shift. In Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, pages 474–484, New Orleans, Louisiana.
  • Schlechtweg (2023) Dominik Schlechtweg. 2023. Human and Computational Measurement of Lexical Semantic Change. Stuttgart, Germany.
  • Schlechtweg et al. (2024a) Dominik Schlechtweg, Pierluigi Cassotti, Bill Noble, David Alfter, Sabine Schulte im Walde, and Nina Tahmasebi. 2024a. More DWUGs: Extending and evaluating word usage graph datasets in multiple languages. In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing, pages 14379–14393, Miami, Florida, USA. Association for Computational Linguistics.
  • Schlechtweg et al. (2025) Dominik Schlechtweg, Tejaswi Choppa, Wei Zhao, and Michael Roth. 2025. CoMeDi shared task: Median judgment classification & mean disagreement ranking with ordinal word-in-context judgments. In Proceedings of Context and Meaning: Navigating Disagreements in NLP Annotation, pages 33–47, Abu Dhabi, UAE. International Committee on Computational Linguistics.
  • Schlechtweg et al. (2019) Dominik Schlechtweg, Anna Hätty, Marco del Tredici, and Sabine Schulte im Walde. 2019. A Wind of Change: Detecting and Evaluating Lexical Semantic Change across Times and Domains. In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, pages 732–746, Florence, Italy. Association for Computational Linguistics.
  • Schlechtweg et al. (2020) Dominik Schlechtweg, Barbara McGillivray, Simon Hengchen, Haim Dubossarsky, and Nina Tahmasebi. 2020. SemEval-2020 Task 1: Unsupervised Lexical Semantic Change Detection. In Proceedings of the 14th International Workshop on Semantic Evaluation, Barcelona, Spain. Association for Computational Linguistics.
  • Schlechtweg and Schulte im Walde (2020) Dominik Schlechtweg and Sabine Schulte im Walde. 2020. Simulating Lexical Semantic Change from Sense-Annotated Data. In The Evolution of Language: Proceedings of the 13th International Conference (EvoLang13).
  • Schlechtweg et al. (2018) Dominik Schlechtweg, Sabine Schulte im Walde, and Stefanie Eckmann. 2018. Diachronic Usage Relatedness (DURel): A framework for the annotation of lexical semantic change. In Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, pages 169–174, New Orleans, Louisiana.
  • Schlechtweg et al. (2021) Dominik Schlechtweg, Nina Tahmasebi, Simon Hengchen, Haim Dubossarsky, and Barbara McGillivray. 2021. DWUG: A large Resource of Diachronic Word Usage Graphs in Four Languages. In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing, pages 7079–7091, Online and Punta Cana, Dominican Republic. Association for Computational Linguistics.
  • Schlechtweg et al. (2024b) Dominik Schlechtweg, Shafqat Mumtaz Virk, Pauline Sander, Emma Sköldberg, Lukas Theuer Linke, Tuo Zhang, Nina Tahmasebi, Jonas Kuhn, and Sabine Schulte im Walde. 2024b. The DURel annotation tool: Human and computational measurement of semantic proximity, sense clusters and semantic change. In Proceedings of the 18th Conference of the European Chapter of the Association for Computational Linguistics: System Demonstrations, pages 137–149, St. Julians, Malta. Association for Computational Linguistics.
  • Schlechtweg et al. (2024c) Dominik Schlechtweg, Frank D. Zamora-Reina, Felipe Bravo-Marquez, and Nikolay Arefyev. 2024c. Sense through time: Diachronic word sense annotations for word sense induction and lexical semantic change detection. Language Resources and Evaluation.
  • Schütze (1998) Hinrich Schütze. 1998. Automatic word sense discrimination. Computational Linguistics, 24(1):97–123.
  • Trott and Bergen (2021) Sean Trott and Benjamin Bergen. 2021. RAW-C: Relatedness of ambiguous words in context (a new lexical resource for English). In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers), pages 7077–7087, Online. Association for Computational Linguistics.
  • Yadav and Schlechtweg (2025) Sachin Yadav and Dominik Schlechtweg. 2025. XL-DURel: Finetuning sentence transformers for ordinal Word-in-Context classification.
  • Zamora-Reina et al. (2022) Frank D. Zamora-Reina, Felipe Bravo-Marquez, and Dominik Schlechtweg. 2022. LSCDiscovery: A shared task on semantic change discovery and detection in Spanish. In Proceedings of the 3rd International Workshop on Computational Approaches to Historical Language Change, Dublin, Ireland. Association for Computational Linguistics.
  • Zamora-Reina et al. (2025) Frank D. Zamora-Reina, Felipe Bravo-Marquez, Dominik Schlechtweg, and Nikolay Arefyev. 2025. Can large language models compete with specialized models in lexical semantic change detection? In Proceedings of the European Conference on Artificial Intelligence (ECAI).